If the circle passing through the point (-1,0) touches the y-axis at (0,2) then what is the equation of the circle?


1
Expert's answer
2021-10-06T13:24:59-0400

Since the required circle touches y-axis at (0,2) and let r be the radius of the circle.

So, the center will be (−r,2).


Equation of circle is

(x+r)2+(y−2)2=r2


Since, it passes through (−1,0),

⇒(r−1)2+4=r2

⇒r=5/2

So, the equation of the circle:

(x+5/2)2+(y−2)2=25/4